A software designer is mapping the streets for a new racing game. All of the streets are depicted as either perpendicular or parallel lines. The equation of the lane passing through A and B is -7x + 3y = -21.5. What is the equation of the central street PQ?


1
Expert's answer
2021-06-23T14:14:38-0400
"-7x+3y=-21.5"

"y=\dfrac{7}{3}x-\dfrac{21.5}{3}"

The slope of the parallel line is "\dfrac{7}{3}."

The slope of the perpendicular line is "-\dfrac{3}{7}."

A.


"-3x+4y=3"

"y=\dfrac{3}{4}x+\dfrac{3}{4}"


B.

"-1.5x-3.5y=-31.5"

"y=-\dfrac{3}{7}x+9"

C.


"2x+y=20"

"y=-2x+20"

D.



"-2.25x+y=-9.75"

"y=\dfrac{9}{4}x-\dfrac{39}{4}"


Answer

B.


"-1.5x-3.5y=-31.5"